# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

### v1

@@ V-B Attack Effectiveness
Takeaway 1: Inversion attack is mostly effective against the KV-cache of the first MHA layer.
Results of KV-cache Collision Attack. For the collision attack, we measure the distance between a locally generated KV-cache and the leaked KV-cache using the Frobenius norm. We experimentally observed that the distances ($dis_{other}$) between the leaked KV-cache and the caches generated from incorrect tokens in the vocabulary approximate a Gaussian distribution, as shown in Figure 8. Consequently, we leverage the statistical properties of this distribution (i.e., its mean and standard deviation) to identify the correct token, whose corresponding distance ($dis_{target}$) will manifest as a statistical outlier.
After experimental analysis, we defined outliers as values falling below 3 standard deviations from the mean of the $dis_{batch}$ distribution. The batch size for local analysis was set to 256. Details of the ablation study can be found in Appendix -C1. Under these settings, the experimental results for the collision attack are presented in the “Collision” column of Table I, demonstrating high reconstruction accuracy across all layers of all models.
We also experimented with truncating the probability-sorted vocabulary on the LLaMA-3.2-1B model, aiming to reduce the time lost on unlikely tokens. The results are presented in Figure 6. We found that even when searching only the top $1/8$ of tokens with the highest predicted probabilities, the average input reconstruction fidelity remains at approximately 96.1% of what is achieved by searching the entire vocabulary. Meanwhile, the attack time is reduced to less than 42.9% of the full search, with the average time to reconstruct a user input from a single layer’s KV-cache dropping from 5.06s to 2.17s. This confirms that truncating the vocabulary based on model-predicted probabilities is a highly effective strategy for accelerating the collision attack.
Figure 6: The effect of truncating the probability-sorted vocabulary on reconstruction fidelity and attack time. Experiments were run with a batch size of 256 and an outlier threshold of $3\sigma_{other}$.
The experimental results for the collision attack demonstrate that this method overcomes the limitations of the inversion attack, which is restricted to specific layers (i.e., the first layer) and attention mechanisms (i.e., MHA). For open-source models, the collision attack can reconstruct user inputs with high accuracy and acceptable efficiency from the KV-cache of any layer. Crucially, it is also effective against the LLaMA-3.1-8B-Distilled model—a model fine-tuned from Meta-LLaMA-3.1-8B—by using the public weights of its open-source base model.
Takeaway 2: Collision attack achieves relatively high accuracy, demonstrating its universality and the high-risk threat it poses to real-world systems.
Collision Attack Enhanced with Prior Knowledge (Collision+). As shown in Table I, our enhanced collision attack (assuming rank $r=8$) is highly potent, achieving near-perfect reconstruction accuracy (approaching 100%) on any layer of any tested open-source model, while also demonstrating notable effectiveness against fine-tuned models.
We also conducted experiments on the LLaMA-3.1-8B-Distilled model using optimal thresholds derived for different assumed ranks ($r$). The results are shown in Figure 7. For the first layer, the actual average rank of the target token is approximately 4; using the optimal threshold for this rank increases the reconstruction accuracy to 138.6% of the unenhanced attack. For the middle layer, the actual rank appears to be greater than 128, as none of the tested enhanced thresholds surpassed the accuracy of the baseline attack. For the last layer, the optimal rank is between 64 and 128, and applying the corresponding enhanced threshold boosts the reconstruction accuracy to approximately 139.0% of the baseline.
@@
Figure 8: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) from a collision attack. The input is an excerpt from “The Bitter Lesson”. The attack uses the LLaMA-3.1-8B model’s weights against the last-layer KV-cache generated by itself (left) and by LLaMA-3.1-8B-Distilled (right). The green dotted line shows the baseline heuristic threshold ($3\sigma_{other}$), while the red dashed line shows the enhanced threshold derived from prior knowledge (assuming rank $r=64$).
To further illustrate this, we analyzed the distance distributions using an excerpt from “The Bitter Lesson” as input. Figure 8 shows these distributions for an attack on the last-layer KV-cache. We observe that for the open-source model (left panel), the baseline heuristic threshold of $3\sigma_{other}$ results in a 0.13% false-positive rate (i.e., 0.13% of $dis_{other}$ distances are misclassified). If we assume the correct token is at rank $r=64$ in the sorted vocabulary, this yields a per-token attack success rate of 91.84%. In contrast, by leveraging a more precise threshold derived from statistical analysis, this success rate can be boosted to a perfect 100%. For the fine-tuned model (right panel), the enhanced threshold improves the per-token success rate from 91.79% to 97.82%. This substantial improvement in per-token accuracy leads to a dramatic increase in the overall reconstruction fidelity for the entire input sequence.
Results of KV-cache Injection Attack. The success of this attack hinges on two factors: the model’s ability to effectively parse the semantic content of the KV-cache and its comprehension of the injected instruction. For our evaluation, we employed the most effective instruction identified in our ablation study, “Repeat the previous content.” (the selection process is detailed in Appendix -C2), to attack the KV-cache data from each target model. As shown in Figure I, the injection attack yielded an average BERTScore of 0.58 and a ROUGE-L score of 0.42 for the reconstructed user inputs. These results are consistently lower than those achieved by the more direct collision attack.
The attack performed best on the LLaMA-7B. This supports the hypothesis that the MHA mechanism preserves more contextual information than the more compressed attention variants (e.g., GQA) within the KV-cache.
Takeaway 4: Absent verbatim recovery, injection attacks can still exfiltrate the core information of user inputs through semantic-level interaction.
The reconstruction accuracy of the injection attack is generally lower than that of the collision attack, and it requires access to the complete KV-cache from all layers. However, its key advantage is speed: the attack’s execution time is equivalent to a single inference step, making it significantly faster than the iterative collision attack. Moreover, the existence of this attack vector imposes an additional requirement on any defense mechanism: it must ensure that a protected KV-cache is rendered unintelligible even to a standard, unprotected LLM.
V-C Evaluation of KV-Cloak
V-C1 Experimental Settings
In order to comprehensively evaluate our proposed KV-Cloak scheme, this section presents a systematic comparison against two baselines: (1) a standard, unprotected system (Plaintext) and (2) a defense based on differential privacy with Gaussian noise (DP). Our evaluation is structured around three core dimensions: Security, Model Accuracy, and Performance Overhead. The results demonstrate that KV-Cloak is the only solution that provides robust security protection with acceptable performance overhead, all without sacrificing model accuracy.
In our KV-Cloak, the obfuscation involves matrix multiplication of KV-cache blocks with invertible secret matrices $S$ and $M$, along with the addition of a random mask $A$.
  • Block Size: To ensure compatibility with the PagedAttention mechanism used in modern inference engines like vLLM, we experimented with common block sizes: 16, 32, and 64.
  • Invertible Matrices $S,M$: To avoid precision loss from matrix inversion during de-obfuscation, we sample $S$ and $M$ from the space of orthogonal matrices.
  • Additive Mask $A$ and Padding: The values in the mask $A$ and the padding for incomplete blocks are kept small to minimize their impact on numerical precision. We sample the elements of $A$ from the range $3\theta_{K},4\theta_{K}]$ and use a value of $1.5\theta_{K}$ for padding. Here, $\theta_{K}$ is the maximum absolute value observed in the K-cache elements when inferring on an excerpt from “The Bitter Lesson”. The same methodology is applied for the V-cache.
To establish a robust DP baseline, we conducted an ablation study (detailed in Appendix -D) to select a configuration that balances utility and privacy for comparison against KV-Cloak. Based on our findings, we selected a practical configuration: the clipping threshold was set to the 50th percentile of the L2 norm distribution observed across the dataset. Subsequently, we applied Gaussian noise calibrated to satisfy ($\epsilon=10^{8},\delta=10^{-5}$)-DP independently to both the K-cache and V-cache.
V-C2 Evaluation of Security

### v3

@@ IV-B2 Results of KV-cache Inversion Attack
Takeaway 1: The Inversion Attack is effective against the first-layer KV-cache of MHA models, with exceptions for architectures possessing high-rank projection matrices.
IV-B3 Results of KV-cache Collision Attack
@@
Figure 3: Distance distributions of target tokens $d_{\text{target}}$ (orange) versus incorrect tokens $d_{\text{other}}$ (blue) in the Collision Attack. The input is an excerpt from “The Bitter Lesson”(see Appendix -B). The attack targets the last-layer KV-cache of LLaMA-3.1-8B (left) and LLaMA-3.1-8B-Distilled (right) using weights from the base model. Vertical lines indicate the heuristic threshold ($3\sigma_{\text{other}}$, green dotted) and the prior-knowledge enhanced threshold ($r=64$, red dashed).
Figure 4: The effect of truncating the probability-sorted vocabulary on reconstruction fidelity and attack time. Experiments were run with a batch size of 256 and an outlier threshold of $3\sigma_{\text{other}}$.
Unlike the Inversion Attack, the Collision Attack relies on forward-pass matching using the Frobenius norm. As shown in Figure 3, we observe that distances for incorrect tokens ($d_{\text{other}}$) approximate a Gaussian distribution, whereas the target token ($d_{\text{target}}$) manifests as a distinct statistical outlier. By setting an outlier threshold of $3\sigma$ below the batch mean (batch size=256, see Appendix -C1), we achieve high reconstruction accuracy across all layers of all tested models (Table I). Crucially, this method overcomes the architectural constraints of MHA; it is effective even against the fine-tuned LLaMA-3.1-8B-Distilled using only public base model weights, demonstrating robustness against parameter divergence.
Efficiency via Probability-Guided Pruning. To enhance practicality, we evaluate truncating the search space based on model-predicted probabilities. Experimental results on LLaMA-3.2-1B (Figure 4) indicate that searching only the top $1/8$ of tokens retains 96.1% of the full-search fidelity. This optimization reduces the average reconstruction time per layer from 5.06s to 2.17s ($43\%$ of the original latency), rendering the attack highly efficient for real-time scenarios.
Takeaway 2: The Collision Attack is a universal threat, achieving high fidelity across diverse architectures (including fine-tuned models) and layers, with high efficiency via probability pruning.
IV-B4 Collision Attack Enhanced with Prior Knowledge (Collision+)
We further evaluate the potency of integrating adversarial prior knowledge (i.e., adaptive thresholding based on assumed token rank $r$).
Efficacy on Fine-tuned Models. As detailed in Table I, assuming a rank $r=8$ yields near-perfect reconstruction ($\approx 100\%$) across all open-source models. For the fine-tuned LLaMA-3.1-8B-Distilled, we analyze the impact of rank assumptions in Figure 5. In the first layer, where the actual average target rank is $\approx 4$, using the optimal threshold boosts accuracy to 138.6% of the baseline. Conversely, the middle layer shows no improvement, implying an actual rank >128>128. In the last layer, optimal rank alignment ($64\leq r\leq 128$) increases accuracy to 139.0% relative to the baseline.
Statistical Justification. To further illustrate this, we analyze distance distributions using “The Bitter Lesson” excerpt (Figure 3). For the open-source model, the heuristic $3\sigma_{\text{other}}$ threshold yields a 0.13% false-positive rate, capping per-token success at 91.84% (assuming $r=64$). By switching to a statistically derived threshold, we eliminate false positives, achieving 100% success. Similarly, for the fine-tuned model, the enhanced threshold improves per-token success from 91.79% to 97.82%, dramatically increasing full-sequence reconstruction fidelity.
IV-B5 Results of KV-cache Injection Attack
We evaluate semantic exfiltration using the optimal directive identified in Appendix -C2: “Repeat the previous content”. As detailed in Table I, this vector achieves an average BERTScore of 0.58 and ROUGE-L of 0.42. While these metrics are lower than those of the exact-match Collision Attack, they confirm significant leakage of the input’s core meaning.
Notably, the attack exhibits peak performance on LLaMA-7B. This supports the hypothesis that the MHA mechanism retains higher contextual fidelity within the KV-cache compared to compressed variants like GQA. Consequently, MHA caches are more semantically intelligible to the model, increasing vulnerability to instruction-based extraction. This imposes a critical requirement for defenses: a protected KV-cache must be rendered semantically unparsable to the LLM itself, preventing the model from being weaponized to interpret stolen contexts.
Takeaway 4: Even without verbatim recovery, Injection Attacks successfully exfiltrate core user intent, necessitating defenses that neutralize semantic intelligibility.
IV-C Attack Robustness under Partial Knowledge Scenarios
To assess practical applicability, we evaluate the robustness of our attacks under a relaxed threat model characterized by two realistic constraints: incomplete data interception and model parameter mismatch.
Incomplete KV-cache Data. Our analysis reveals distinct data requirements across attack vectors. The Injection Attack necessitates the full, cross-layer context and fails when data is fragmented. Conversely, the Collision Attack exhibits superior robustness; possessing the KV-cache from any single layer is sufficient to perform high-fidelity input reconstruction.
Model Mismatch. We further evaluate the scenario where the adversary’s local model differs from the target service model. Experiments confirm that attack effectiveness hinges on the correlation of weight parameters.
  • Fine-tuning Mismatch (Gray-box): We targeted a fine-tuned model (LLaMA-3.1-8B-Distilled) assuming the adversary only possesses the open-source base model (LLaMA-3.1-8B). Results in Table I show that the attack remains highly effective, suggesting that fine-tuning preserves the fundamental weight correlations exploited by our mechanism.
  • Architectural Mismatch (Black-box): In contrast, cross-architecture attacks (e.g., using LLaMA-7B to attack LLaMA-2-7B) fail completely. Reconstruction similarity drops to levels statistically indistinguishable from random guessing (<0.1<0.1). This confirms that the Collision Attack relies on the algebraic correspondence of weight parameters, which is absent across disjoint architectures.
V KV-Cloak: A Lightweight KV-cache Defense
V-A Motivation for KV-Cloak
Limitations of Existing Privacy-Preserving Techniques. Current privacy-preserving techniques fundamentally struggle to balance security, efficiency, and utility when applied to the massive, latency-sensitive KV-cache. We analyze their specific limitations below:
  • Cryptographic Methods (Overhead Bottleneck): Standard cryptographic techniques [4, 34], such as symmetric AES encryption, provide confidentiality by ensuring data remains in ciphertext during storage and transmission, requiring decryption only for legitimate use. A more advanced approach, Homomorphic Encryption (HE) [26, 2], allows for direct computation on encrypted data, theoretically enabling parts of the attention mechanism to operate on the KV-cache without ever decrypting it. However, despite their strong security guarantees, the immense computational overhead and latency introduced by these methods are prohibitive. Given that the KV-cache can be tens or hundreds of gigabytes, applying full encryption or HE is unsuitable for the high-throughput, low-latency requirements of LLM inference.
  • Differential Privacy (Utility Trade-off): Applying Differential Privacy involves injecting calibrated noise into the Key-Value vectors to mask individual data points [1, 20, 42]. However, the sparse and sensitive nature of attention mechanisms creates a severe utility-privacy trade-off. To achieve a meaningful level of privacy, the required amount of noise would significantly degrade the LLM’s inference accuracy to an unacceptable degree.
  • KV-Shield (Security and Compatibility Flaws): KV-Shield [41] is the only existing lightweight obfuscation scheme for KV-cache. It synchronously permutes attention weight rows ($W_{q},W_{k},W_{v}$) to shuffle cache elements, obfuscating them before an attention score is calculated. The obfuscated attention output is then de-obfuscated for subsequent steps. We identify two critical failures in this design: (1) Security Flaws. First, the simple shuffling preserves the inherent statistical distribution of the data, leaving it highly vulnerable to our proposed Collision Attacks. Second, the reliance on a fixed obfuscation key lacks dynamic randomness, rendering the scheme defenseless against CPA. (2) Architectural Incompatibility. The shuffling disrupts relative positioning, making it incompatible with RoPE used in SOTA models (e.g., LLaMA, Qwen).
Design Objectives. The deficiencies of prior works necessitate a specialized defense for KV-cache. We define three mandatory design goals for a practical solution:
  • Robust Security: The mechanism must resist targeted reconstruction attacks by effectively masking both the algebraic and statistical properties of the cache.
  • Lossless Model Fidelity: It must preserve the exact mathematical equivalence of the attention mechanism, ensuring zero degradation in generation quality.
  • Negligible Overhead: The defense must operate with minimal latency, ideally shifting computational costs offline to maintain high inference throughput.
To meet these objectives, we propose KV-Cloak, a novel defense mechanism that synergizes reversible matrix obfuscation with operator fusion to deliver robust security, lossless model fidelity, and negligible inference overhead.
V-B Naive Defense: Reversible Linear Obfuscation
To mitigate privacy risks with minimal overhead, we first consider a naive defense that employs reversible linear transformations to obscure the statistical properties of the KV-cache. In the context of modern inference frameworks (e.g., vLLM), we denote the key vectors for a single attention head within a PagedAttention block as a matrix $K\in\mathbb{R}^{b\times d}$, where $b$ is the block size and $d$ is the head dimension. Since the protection mechanism applies analogously to value vectors, we focus our analysis on $K$. The obfuscation transformation is defined as:
$K^{\prime}=SKM,$ (7)
where $S\in\mathbb{R}^{b\times b}$ and $M\in\mathbb{R}^{d\times d}$ are secret, randomly generated invertible matrices. This transformation aims to conceal the raw content of $K$ while preserving the matrix dimensions necessary for storage allocation.
Security Analysis. Despite its operational simplicity, this scheme is fundamentally insecure under a CPA model due to its fixed linear structure. While a real-world adversary cannot generate arbitrary matrix $K$ directly (as $K$ is constrained by the model’s embedding projection of valid tokens), we demonstrate that this constraint does not preclude a full compromise. An adversary can circumvent this restriction by mounting a differential attack through carefully crafted prompts. Specifically, the adversary chooses two inputs yielding plaintexts $K_{1}$ and $K_{2}$, thereby controlling the difference $\Delta K=K_{1}-K_{2}$. Due to the linearity of the scheme, the observed ciphertext difference is $\Delta K^{\prime}=S(\Delta K)M$. By systematically crafting inputs such that $\Delta K$ approaches a series of standard basis matrices (i.e., matrices with a single non-zero entry), the adversary can isolate and solve for the columns of $S$ and the rows of $M$. This algebraic attack enables full recovery of the secret matrices (up to a scalar ambiguity) with a computational complexity of only $O(b^{2}d+bd^{2})$, rendering the naive defense ineffective against determined adversaries.
V-C One-Time Pad Block-wise Shuffling
Figure 6: Illustration of the KV-Cloak block-wise shuffling mechanism. By permuting $K$ and $V$ vectors within a block, the mechanism eliminates positional side channels while maintaining mathematical equivalence in the attention output.
V-C1 Eliminating Redundant Positional Information in KV-cache
To enhance the security of the obfuscation scheme, we introduce additional randomness by eliminating the architectural redundancy of the KV-cache. We observe that the physical memory ordering of $k,v$ vectors is computationally superfluous during inference, as positional semantics are intrinsic to the vectors via RoPE.
Mechanism. We implement a dynamic, block-wise random permutation. By reordering $k,v$ pairs within each block while maintaining their internal correspondence, we decorrelate the physical storage index from the logical token sequence. As illustrated in Figure 6, this operation preserves the mathematical invariance of the attention mechanism while imposing a combinatorial complexity barrier of $b!$ on the adversary. Given typical block sizes $b\in\{16,32,64\}$ [17], this renders brute-force matching computationally infeasible.
Formalization. The enhanced obfuscation employs a locally generated, one-time pad permutation matrix $\hat{P}$ before applying the linear transformations:
$K^{\prime}=S\hat{P}KM.$ (8)
Crucially, for efficiency, $\hat{P}$ is ephemeral and does not require storage. The inference engine performs subsequent computations directly on the de-obfuscated permuted state $\hat{P}K$ without reconstructing the original order, thereby incurring zero additional storage overhead for the permutation key.
V-C2 Rank Preservation and Implicit Key Recovery
Figure 7: Value distribution of KV-cache elements ($K$ and $V$) for LLaMA-3.2-1B. The data is collected during inference on “The Bitter Lesson”.
A critical challenge arises when the original matrix $K$ exhibits low rank (e.g., rank $1$ in cases of identical repeated tokens). In such scenarios, the permutation entropy collapses, as row shuffling becomes statistically invisible. To guarantee obfuscation robustness, we introduce a secret additive mask matrix $A$ before permutation. Thereby, we can ensure the transformed matrix $(K+A)$ maintains sufficient rank to preserve cryptographic strength of $\hat{P}$. However, explicitly storing the unique one-time pad $\hat{P}$ for every block would reintroduce significant storage overhead. We resolve this by designing $A$ to enable implicit key recovery without $\hat{P}$.
Magnitude-based Positional Embedding. We exploit the numerical characteristics of attention activations. As shown in Figure 7, elements of $k,v$ vectors are statistically bounded by a threshold $\theta_{K}$ (e.g., typically <100<100). Leveraging this, we construct $A$ to function as a “positional beacon” by embedding values significantly larger than $\theta_{K}$ into specific coordinates. This ensures that the matrix $(K+A)$ maintains enough rank.
Zero-Storage Recovery. Since the “beacons” in $A$ dominate the magnitude of $K$, the permuted mask term $\hat{P}A$ remains statistically separable from $\hat{P}K$ within the ciphertext $\hat{P}(K+A)$. During de-obfuscation, the system dynamically identifies the row permutation indices by detecting these high-magnitude outliers, allowing it to reconstruct $\hat{P}A$ on-the-fly. Subtracting this reconstructed mask yields the permuted state $\hat{P}K$ required for inference. This mechanism effectively offloads the storage of $\hat{P}$ into the data itself.
Formalization. The complete obfuscation logic integrates this additive masking with the linear and permutation layers:
$K^{\prime}=S\hat{P}(K+A)M,$ (9)
[…]
V-C3 Security Analysis
The security of KV-Cloak relies fundamentally on the confidentiality of its secret matrices ($S,M,A$). Under a standard CPA model, an adversary attempting to recover these keys faces a prohibitive computational barrier. Our analysis indicates that the time complexity to solve for the keys is $O((b^{2}d+bd^{2})\cdot b!)$. Crucially, the factorial term $b!$, introduced by the dynamic block-wise permutation, renders brute-force key recovery computationally infeasible.
This computational hardness is complemented by layered security mechanisms designed to mitigate specific attack vectors: (1) The one-time pad permutation severs the explicit token-to-vector positional correspondence at an information-theoretic level, invalidating the premise of Collision Attacks. (2) A reversible algebraic transformation via secret matrices ($S,M$) completely disrupts the cache’s statistical distribution properties (e.g., mean, variance), neutralizing profiling and data mining attacks. (3) The combined obfuscation renders the cache semantically unintelligible and unparsable by pristine (unmodified) models, thereby thwarting Injection Attacks that attempt to leverage the model’s own capabilities.
V-D Implicit Obfuscation via Operator Fusion
To minimize the runtime latency overhead introduced by the obfuscation and de-obfuscation processes, we propose an operator fusion strategy. This approach integrates the obfuscation operators directly into the model’s weight matrices offline, thereby rendering the online computational impact negligible while maintaining strict mathematical equivalence.
Transformed Vectors and Invariance. We introduce two secret, invertible matrices, $M_{1},M_{2}\in\mathbb{R}^{d\times d}$, to transform the query, key, and value vectors ($q,k,v$) and the output projection $W_{o}$. The transformed vectors ($q^{m},k^{m},v^{m}$) and the modified output weight $W_{o}^{m}$ are defined as:
This design ensures that the core attention mechanism remains invariant. Specifically, the attention scores (scaled dot-product of query and key) are preserved:
$=q_{i}(M_{1}M_{1}^{-1})^{\top}k_{j}^{\top}=q_{i}k_{j}^{\top}.$
Similarly, the output of the attention head is mathematically identical to the unprotected, guaranteeing lossless accuracy:
$=v_{j}M_{2}(M_{2}^{-1})W_{o}^{\top}=v_{j}W_{o}^{\top}.$
Fusion under RoPE. Implementing this fusion requires incorporating the input projection $x$ and the position-dependent RoPE matrix $R_{\Theta,i}^{d}$. The transformed vectors are expressed as:
From Eq. 13, fusing $M_{2}$ into the value and output weights is straightforward: we pre-compute $(W_{v}^{m})^{\top}=W_{v}^{\top}M_{2}$ and $W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}$. However, fusing $M_{1}$ into $W_{k}$ and $W_{q}$ presents a challenge due to the intermediate application of RoPE. The fusion is algebraically feasible if and only if the secret matrix $M_{1}$ and the RoPE matrix $R_{\Theta,i}^{d}$ commute, i.e., $R_{\Theta,i}^{d}M_{1}=M_{1}R_{\Theta,i}^{d}$. In this case, Eq. 13 simplifies to:
$k^{m}=xW_{k}^{\top}(M_{1}R_{\Theta,i}^{d})=x(W_{k}^{\top}M_{1})R_{\Theta,i}^{d}.$ (14)
This reordering allows us to absorb $M_{1}$ into the weight matrix as $(W_{k}^{m})^{\top}=W_{k}^{\top}M_{1}$. A symmetric logic applies to $W_{q}$.
To satisfy this commutativity constraint, we structurally design $M_{1}$ as a invertible block-diagonal matrix composed of $2\times 2$ rotation-scaling sub-matrices, analogous to the structure of RoPE (detailed in Appendix -A). Additionally, we constrain $M_{2}$ to be a random invertible rotation-scaling matrix and calibrate the scaling factors of both matrices. This ensures that after the transformation with mask $A$, the padding tokens remain statistically distinguishable as outliers for implicit key recovery.
Offline Computation and Online Efficiency. Consequently, the obfuscation operators are fully fused into the attention layer parameters offline. The new pre-computed weights are:
During online inference, the KV-Cloak obfuscation is applied to the cache generated by these fused weights:
$K^{\prime}=S\hat{P}(K^{m}+A),$ (16)
By eliminating explicit online matrix multiplications with $M$, the cost for protecting one KV-cache block is reduced to the multiplications with $S$, $\hat{P}$, and $S^{-1}$. This totals approximately $b^{3}+2b^{2}d$ floating-point operations. For an LLaMA-3.1-8B instance ($b=16,d=128,D=4096$), this overhead represents merely 0.83% of the KV-cache re-computation cost (detailed in Appendix -F), verifying that KV-Cloak achieves robust security with minimal performance impact.